OpenStudy (zepp):

I have a tricky question with 3 variables but only with 2 equations :| \(a+b+c=45\) \(59a+41b+40c=2008\) P.S.: I only want hints, thank you all :)

OpenStudy (anonymous):

a b c z ... that's 4 lol :A

OpenStudy (zepp):

Fixed D; It was a c, not a z >.>

OpenStudy (anonymous):

Free variable...

OpenStudy (anonymous):

Hint: There will not be a unique solution.

OpenStudy (zepp):

Trial and error? D:

OpenStudy (anonymous):

What context is this problem in? The way I would usually solve this uses methods from linear algebra.

OpenStudy (anonymous):

You use the same method of matrices but now you have a free variable, meaning your last variable is free to be whatever as long as it makes sense

OpenStudy (zepp):

Context? I'm still in 10th grade :( but I learned a bit of linear algebra.

OpenStudy (anonymous):

By context I just mean what you're learning right now. As I'm sure you've noticed, the methods used to solve a problem tend to change a lot as you progress through learning math :)

OpenStudy (zepp):

Oh uh, solving equation systems with substitution, reduction and comparison, you know, basic stuffs. :)

OpenStudy (kira_yamato):

You can set solution as parametric equations, like in this case,

OpenStudy (kira_yamato):

Then depending on question requirement, you put in a certain value for the parameter.

OpenStudy (anonymous):

The way we do it in linear algebra is to find the null space of the matrix that represents the system of equations, and then find a particular solution, and the solution space is the sum of a the particular solution and the null space. You can do basically the same thing without any of that terminology or matrix-oriented methodology, basically using elimination. First step, you can find a particular solution by using elimination the same way you normally would.

OpenStudy (zepp):

After some substitutions and eliminations I got some interesting things \(c=163+18a\) \(95a + 5b=388\) \(59a + 5b +2c= 714\)

OpenStudy (zepp):

And now the problem is, if I substitute the first one into \(a+b+c=45\) I would get a negative number, and from the question, a,b,x are > 0

OpenStudy (zepp):

Arrgh a,b,c*

OpenStudy (anonymous):

Aha, I didn't know that the three variables had to be positive. That ends up not being a problem at all, because there is actually an infinite line of solutions. Once you have a particular solution, you can modify it to make it so that all three variables are positive.

OpenStudy (anonymous):

Here's a hint to make this easier on you: the quickest way to find a particular solution is almost always to set one variable equal to zero, and then solve for the other two.

OpenStudy (kira_yamato):

You can use parameters.

OpenStudy (zepp):

@nbouscal Do you mean something like stuffswithvariable = 0?

OpenStudy (anonymous):

I just mean pick a, b, or c, and set it equal to zero. Then solve the resulting system of equations. So for example, if you set a equal to zero, you will be left with b+c=45 and 41b+40c=2008

OpenStudy (zepp):

Oh I see, let me try.

OpenStudy (kira_yamato):

For me, I tend to write it in parameters first, then plug in suitable values for the parameters. \[a = 9\frac{1}{18}+\frac{1}{18}t\] \[b = 35\frac{17}{18}+1\frac{1}{18}t\] \[c = t\] For\[t \in \mathbb{R}\]

OpenStudy (zepp):

Okay, by setting up 1 variable to 0, I get If a = 0 b+c=45 41b+40c=2008 So a=0 b=208 c=-163 Other give decimal numbers D:

OpenStudy (zepp):

@Kira_Yamato I didn't learn parametrization.

OpenStudy (kira_yamato):

Souka... I'm sorry...

OpenStudy (anonymous):

Yeah, I did it with a=0 also. Anyway, now you have a particular solution to the system. The next part is to learn about the null space (in linear algebra terms). Basically it provides you with a set of numbers that you can add to your solutions to get another solution. The way you find those numbers is by setting the equations equal to zero, and finding a solution. So you have a+b+c=0 and 59a+41b+40c=0. Easiest way to solve them is to use elimination. The intuition on this is actually really cool. If you find a set of numbers that, when plugged into the equations, return zero... you can add that set of numbers to a solution to the original equations and you'll still have a solution, because you're just adding zero.

OpenStudy (anonymous):

I should point out that Kira's method of solving using parametric equations is actually a bit quicker and easier than my method if you haven't yet learned linear algebra. Once you've learned linear algebra I think my method's a little quicker, but it might be personal preference.

OpenStudy (zepp):

Alright now I eliminated stuffs from the equations above, I get 18b+19c=0

OpenStudy (zepp):

Do I have to find a set that could match the thingy above ^

OpenStudy (anonymous):

Don't know if it helps at all http://wiki.answers.com/Q/Is_it_possible_tofind_three_unknowns_with_two_equations

OpenStudy (anonymous):

18b+19c=0 is correct, but isn't all of the information. You also need some information about a. What I had was that -18a+c=0, and that 19a+b=0. With that information we can find the basis of the nullspace. I had it as a=1, b=-19, and c=18. You could get it in other ways and get other numbers, but the numbers will always be multiples of those numbers. We generally represent that as \(N=c\left[ \begin{array}{c} 1\\-19\\18 \end{array}\right]\) where c is any constant. This is called a linear combination (that term makes more sense when the space in question has more than one dimension). So basically your solution space is your particular solution (a=0, b=208, c=-163) plus anything in the null space, which is any constant multiplied by the solutions above. Since we want to find a solution where all three are positive, we can set \(c=10\), so \(a=0+10\cdot1, b=208+10\cdot-19, c=-163+18\cdot10\), or \(a=10, b=18, c=17\). It turns out this is the only solution where all three variables are positive.

OpenStudy (zepp):

How did you find a=1, b=-19, and c=18? D:

OpenStudy (anonymous):

Well, I had the two equations, -18a+c=0 and 19a+b=0. I got those two equations by using elimination. Once I had them, since I've been treating a as my free variable, I just set a=1 and there you go.

OpenStudy (anonymous):

make third equation by solving these two than solve simultaneously

OpenStudy (zepp):

I think I understood pretty much everything here, thanks! :D And just a little question: how to choose the constant c?

OpenStudy (anonymous):

I just made the observation that to bring the negative solution -163 up to a positive one, we would have to have c be at least 10. 9*18=162<163. A nice side effect of learning it this way is that when you do take more linear algebra, all of this will look kind of familiar! :)

OpenStudy (zepp):

Alright, you are awesome! :D Thank you very much!